OpenStudy (anonymous):

Randall borrowed $850 from his parents to make some repairs on his car. He promised to repay the loan by giving his parents at least $94 from his paycheck each week. Assuming Randall sticks to his promise, which of the following is not a possible scenario for the amount of debt Randall has remaining? mEDAL ? HELP ME PLZZZ

OpenStudy (igreen):

This is just like the other one, your equation is going to be 94x (94 dollars a week being paid) = 850(total amount to pay off). What are your choices?

OpenStudy (anonymous):

IS IT After 5 weeks, Randall is $125 in debt.

OpenStudy (igreen):

Let's find out. Use the equation. 94x = 850 Plug in 5 weeks:(x = 5) 94(5) = 850 470 = 850 850 - 470 = 380. No, Randall will be $380 in debt.

OpenStudy (anonymous):

U WRONG THERE'S NO 380 IN THE ANSWER CHOICE

OpenStudy (igreen):

No, I'm not wrong. Just post your answer choices and I'll tell you which one it is.

OpenStudy (anonymous):

After 5 weeks, Randall is $125 in debt. After 6 weeks, Randall is $175 in debt. After 7 weeks, Randall is $225 in debt. After 8 weeks, Randall is $75 in debt.

OpenStudy (igreen):

Okay, let's find out 6 weeks: Use the equation 94x = 850: 94x = 850 Plug in 6 weeks: 94(6) = 850 Multiply: 564 = 850 Subtract: 850 - 564 = 286. So Option B is incorrect because he will be $286 in debt and not $175

OpenStudy (igreen):

Let's find out 7 weeks: 94x = 850 Plug in 7 weeks: 94(7) = 850 Multiply: 658 = 850 Subtract: 850 - 658 = 192. So Option C is incorrect because Randall will be $192 in debt and not $225.

OpenStudy (igreen):

Let's find 8 weeks: 94x = 850 Plug in 8 weeks: 94(8) = 850 Multiply: 752 = 850 Subtract: 850 - 752 = $98.

OpenStudy (igreen):

Ugh, I'm confused all of the Option seem incorrect. @ganeshie8

OpenStudy (anonymous):

:/ IKR @iGreen

OpenStudy (igreen):

Let's hope @ganeshie8 can figure this out for us.

ganeshie8 (ganeshie8):

say his remaining balance is \(\large B\)

ganeshie8 (ganeshie8):

since he is paying *atleast* $94 each week, we get : \[\large B \le 850-94x\]

ganeshie8 (ganeshie8):

after x=5 weeks, his outstanding balance will be : \[\large B \le 850-94*5\] \[\large B \le 370\]

ganeshie8 (ganeshie8):

so first option After 5 weeks, Randall is $125 in debt. is not your answer. similarly check other options

OpenStudy (igreen):

How is it not the answer? It says less than or equal to 370..doesn't that include 125?

ganeshie8 (ganeshie8):

since 125 is less than 370, this is possible... but the question is about whats NOT possible, right ?

OpenStudy (anonymous):

@iGreen giving me wrong answer wrong explANATION I'M GIVING MEDALS FOR NO REASON

ganeshie8 (ganeshie8):

lol no, this is a trick question

ganeshie8 (ganeshie8):

the equation is correct, just check that the given balance in the option is less than the balance given by equaiton

OpenStudy (igreen):

Oh, yeah it says which one is NOT a possible scenario.

OpenStudy (anonymous):

nOT ONLY THAT HE HELPED ME WITH A LOT OF QUESTION THE ANSWERS HE GAVE ARE WRONG AND I NEED MY GRADE UP SO I CAN GRADUATE NOT DOWN SMH